# Loopy: Taming Audio-Driven Portrait Avatar with Long-Term Motion Dependency

### v2

@@ 3.3 Audio Condition Module
For the audio condition, we first use wav2vec (Baevski et al., 2020; Schneider et al., 2019) for audio feature extraction. Following the approach in EMO, we concatenate the hidden states from each layer of the wav2vec network to obtain multi-scale audio features. For each video frame, we concatenate the audio features of the two preceding and two succeeding frames, resulting in a 5-frame audio feature as audio embeddings $c_{a}udio$ for the current frame. Initially, in each residual block, we use cross-attention with noisy latents as the query and audio embedding $c_{audio}$ as the key and value to compute an attended audio feature. This attended audio feature is then added to the noisy latents feature obtained from self-attention, resulting in a new noisy latents feature. This provides a preliminary audio condition.
@@
Additionally, as illustrated in Figure 4, we introduce the Audio-to-Latents module. This module maps conditions that have both strong and weak correlations with portrait motion, such as audio, to a shared motion latents space. These mapped conditions serve as the final condition features, thereby enhancing the modeling of the relationship between audio and portrait movements based on motion latents. Specifically, we maintain a set of learnable embeddings. For each input condition, we map it to a query feature using a fully connected (FC) layer, while the learnable embeddings serve as key and value features for attention computation to obtain a new feature based on the learnable embeddings. These new features, referred to as motion latents, replace the input condition in subsequent computations. These motion latents are then dimensionally transformed via an FC layer and added to the timestep embedding features for subsequent network computation. During training, we sample an input condition for the Audio-to-Latents module with equal probability from audio embeddings and facial movement-related features such as landmarks, face absolute motion variance, and face expression motion variance. During testing, we only input audio to generate motion latents. By leveraging features strongly correlated with portrait movement, the model utilizes learnable embeddings to control motion. Consequently, transforming audio embeddings into motion latents can also allows for more direct influence on portrait motion.
@@ 3.4 Training Strategies
Conditions Mask and Dropout. In the Loopy framework, various conditions are involved, including the reference image $c_{ref}$, audio features $c_{audio}$, preceding frame motion frames $c_{mf}$, and motion latents $c_{ml}$ representing audio and facial motion conditions. Due to the overlapping nature of the information contained within these conditions, to better learn the unique information specific to each condition, we used distinct masking strategies for the conditions during the training process. During training, $c_{audio}$ and motion latents are masked to all-zero features with a 10% probability. For $c_{ref}$ and $c_{mf}$, we design specific dropout and mask strategies due to their conflicting relationship. $c_{mf}$ also provides appearance information and is closer to the current clip compared to $c_{ref}$, leading the model to heavily rely on motion frames rather than the reference image. This can cause color shifts and artifacts in long video sequences during inference. To address this, $c_{ref}$ has a 15% probability of being dropped, meaning the denoising U-Net will not concatenate features from the reference network during self-attention computation. When $c_{ref}$ is dropped, motion frames are also dropped, meaning the denoising U-Net will not concatenate features from the reference network during temporal attention computation. Additionally, motion frames have an independent 40% probability of being masked to all-zero features.
Multistage Training. Following AnimateAnyone (Hu, 2024) and EMO (Tian et al., 2024), we employ a 2-stage training process. In the first stage, the model is trained without temporal layers and the audio condition module. The inputs to the model are the noisy latents of the target single-frame image and reference image latents, focusing the model on an image-level pose variations task. After completing the first stage, we proceed to the second stage, where the model is initialized with the reference network and denoising U-Net from the first stage. We then add the inter-/intra-clip temporal layers and the audio condition module for full training to obtain the final model.
@@
where $e_{audio}$ includes all conditions $c_{ref}$, $c_{audio}$, $c_{mf}$, $c_{ml}$, $e_{ref}$ but mask $c_{audio}$ to all-zero features, and $e_{base}$ masks $c_{audio}$ to all-zero features and removes the concatenation of reference network features in the denoising U-Net’s self-attention. This approach allows us to control the model’s final output in terms of adherence to the reference image and alignment with the audio. The audio ratio is set to 5 and the reference ratio is set to 3. We use DDIM sampling with 25 denoising steps to complete the inference.
@@ 3.5 Experiments
Datasets. For training Data, We collected talking head video data from the internet, excluding videos with low lip sync scores, excessive head movement, extreme rotation, or incomplete head exposure. This resulted in 160 hours of cleaned training data. Additionally, we supplemented the training data with public datasets such as HDTF (Zhang et al., 2021). For the test set, we randomly sampled 100 videos from CelebV-HQ (Zhu et al., 2022) (a public high-quality celebrity video dataset with mixed scenes) and RAVDESS (Kaggle, ) (a public high-definition indoor talking scene dataset with rich emotions). To test the generalization ability of diffusion-based models, we also collected 20 portrait test images, including real people, anime, side face, and humanoid crafts of different materials, along with 20 audio clips, including speeches, singing, rap and emotionally rich speech. We refer to this test set as the openset test set.
@@
Implementation Details. We trained our model using 24 Nvidia A100 GPUs with a batch size of 24, using the AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of 1e-5. The generated video length was set to 12 frames, and the motion frame length was set to 124 frames, representing the preceding 124 frames of the current 12-frame video. After temporal squeezing, this was compressed to 20 motion frame latents. During training, the reference image was randomly selected from a frame within the video clip. For the facial motion information required to train audo-to-motion module, we used DWPose (Yang et al., 2023) to detect facial keypoints for the current 12 frames. The variance of the absolute position of the nose tip across these 12 frames was used as the absolute head motion variance. The variance of the displacement of the upper half of the facial keypoints (37 keypoints) relative to the nose tip across these 12 frames was used as the expression variance. The training videos were uniformly processed at 25 fps and cropped to 512x512 portrait videos.
@@
Metrics. We assess image quality with the IQA metric (Wu et al., 2023), video motion with VBench’s smooth metric (Huang et al., 2024), and audio-visual synchronization using SyncC and SyncD (Prajwal et al., 2020). For the CelebvHQ and RAVDESS test sets, which have corresponding ground truth videos, we will also compute FVD (Unterthiner et al., 2019), E-FID (Tian et al., 2024) and FID metrics for comparison. Additionally, to compare the global motion (denoted as Glo) and dynamic expressions (denoted as Exp) of the portrait, we calculated the variance values based on key points of the nose and the upper face, specifically excluding the mouth area. We also computed the values for the ground truth video as a reference for comparison. For the openset testset, which lacks ground truth video references, we conducted subjective evaluations. Ten experienced users were invited to assess six key dimensions: identity consistency, video synthesis quality, audio-emotion matching, motion diversity, naturalness of motion, and lip-sync accuracy. For each case, participants were required to identify the top-performing method in each dimension.
@@ 3.5.1 Results and Analysis
Performance in Complex Scenarios. CelebV-HQ includes videos of celebrities speaking in various scenarios, including films and interviews, both indoors and outdoors, with diverse portrait poses. This makes testing on this dataset effectively simulate real-world usage conditions. As shown in Table 1, our method significantly outperforms the compared methods in most metrics, as evidenced by the comparison videos provided in the supplementary materials. Regarding motion-related metrics, although not the best, our results are similar to those of the ground truth in terms of smoothness. In the dynamic expression metric (Exp), our method closely matches GT, more so than the compared methods. For global motion (Glo), our performance is similar to that of EchoMimic. However, it is evident that our method has a distinct advantage in video synthesis quality and lip-sync accuracy.
@@
SadTalker 2.953 3.843 8.765 171.848 1746.038 36.648 0.9964 0.554 0.270 2.248
Hallo 3.505 4.130 9.079 53.992 742.974 35.961 0.9946 0.499 0.255 2.426
VExpress 2.946 3.547 9.415 117.868 1356.510 65.098 0.9957 0.020 0.166 2.414
EchoMimic 3.307 3.136 10.378 54.715 828.966 35.373 0.9926 2.259 0.640 3.018
GT - - - - - - 0.9937 3.249 0.506 -
Loopy 3.780 4.849 8.196 49.153 680.634 33.204 0.9949 2.233 0.452 2.307
@@
SadTalker 3.840 4.304 7.621 22.516 487.924 32.343 0.9955 0.604 0.120 3.270
Hallo 4.393 4.062 8.552 38.471 537.478 19.826 0.9942 0.194 0.080 3.785
VExpress 3.690 5.001 7.710 62.388 982.810 26.736 0.9962 0.007 0.039 3.901
EchoMimic 4.504 3.292 9.096 54.115 688.675 21.058 0.9924 0.641 0.184 3.350
GT - - - - - - 0.9917 3.335 0.317 -
Loopy 4.506 4.814 7.798 16.134 394.288 17.017 0.9923 2.962 0.343 3.132
Figure 5: User voting comparisons on the openset test set. The first row includes experimental results for various categories of input images, while the second row includes experimental results for various categories of input audio.
Table 3: Experiments on the effectiveness of the proposed modules.
@@
Full Model 4.507 6.303 7.749 0.9932
w/o inter-clip temp. 4.335 6.104 8.129 0.9942
w/o TSM 4.386 6.054 8.235 0.9922
w/o A2L 4.428 5.999 8.351 0.9922
single temp. + 20 mf 4.072 6.201 8.309 0.9752
$s=1,\ r=2$ 4.461 5.919 8.245 0.9940
$s=2,\ r=2$ 4.453 5.855 8.326 0.9930
$s=3,\ r=2$ 4.443 6.083 8.161 0.9930
$s=4,\ r=1$ 4.424 6.219 8.004 0.9931
mean sample 4.452 5.907 8.199 0.9931
random sample 4.438 6.098 8.144 0.9932
Performance in Emotional Expression. RAVDESS is a high-definition talking scene dataset containing videos with varying emotional intensities. It effectively evaluates the method’s performance in emotional expression. As demonstrated by the E-FID (Tian et al., 2024) metric, our method outperforms the compared methods. This is corroborated in the motion dynamics metrics Glo and Exp, where our results are closer to the ground truth. Although our lip-sync accuracy is slightly inferior to VExpress, it is important to note that the results generated by VExpress generally lack dynamic motion, as indicated by the Glo, Exp, and E-FID metrics. This static nature can provide an advantage when measured with SyncNet for lip-sync accuracy.
Performance in Openset Scenarios. We compared different input styles (real people, anime, humanoid crafts and sife face) and various types of audio (speech, singing, rap, and emotional audio) to evaluate the robustness of the methods. As shown in Table 3, Loopy consistently outperforms the compared methods across these diverse scenarios.
@@ 3.5.2 Ablation Studies
Analysis of Key Components. We analyzed the impact of the two key components of the Loopy, the inter/intra-clip temporal layer and the audio-to-latents module. For the former, we conducted two experiments: (1) removing the dual temporal layer design and retaining a single temporal layer to handle the temporal relationships of both the last and current clips, similar to methods like EMO and Hallo; (2) removing the temporal segment module and retaining the design of 4 motion frames as in other methods. For the latter, we removed the audio-to-latents module and retained only cross-attention for audio feature injection. The results, listed in Table 3, show that the dual temporal layer design improves temporal stability and image quality, and its removal leads to performance degradation. Removing the temporal segment module prevents the model from learning motion style information from long-term motion dependency, resulting in a decline in overall motion quality, including expressiveness and temporal stability. Removing the audio-to-latents module also reduces the overall quality of synthesized visuals and motion. This is because, during the training of the audio-to-latents module, spatial conditions are added for mixed training. Spatial conditions provide clearer motion guidance compared to audio alone, facilitating easier model convergence. These results validate the effectiveness of our method.
Impact of Long-Term Temporal Dependency. We also investigated the impact of long-term motion dependency on the results and list the results in Table 3. Initially, we compared the effects of extending the motion frame length to 20 under a single temporal layer setting. We observed that while this approach enhances the dynamics of the model’s output, it significantly degrades the overall image quality. In contrast, on the full model with 20 motion frames ($s=4,\ r=1$), the addition of inter/intra-clip temporal layers indeed improved the overall results. Regarding the settings of s and r, we experimented with various values. Starting with a fixed $r=2$, we found that smaller values of s resulted in poorer overall performance. We attribute this to the disparity in FPS between the target and motion frames, which complicates inter-clip temporal modeling due to the lower proportion of motion frames. As s increases, the number of motion frames also increases, allowing the inter-clip temporal layer to effectively model cross-clip temporal relationships, thereby significantly enhancing overall performance. We also compared the effects of r with the same $s=4$, where $r=1$. A smaller r indicates a narrower temporal coverage, which slightly reduces performance compared to the full model ($s=4,\ r=2$). Finally, we explored different sampling strategies for motion frames within the temporal segment module, comparing methods such as average pooling and random single-frame sampling. The uniform sampling approach of the full model proved more effective, likely because it provides more stable interval information, which benefits the inter-clip temporal layer in learning long-term motion information.

### v3

@@ 3.3 Audio Condition Module
For the audio condition, we first use wav2vec (Baevski et al., 2020; Schneider et al., 2019) for audio feature extraction. Following the approach in EMO (Tian et al., 2024), we concatenate the hidden states from each layer of the wav2vec network to obtain multi-scale audio features. For each video frame, we concatenate the audio features of the two preceding and two succeeding frames, resulting in a 5-frame audio feature as audio embeddings $c_{audio}$ for the current frame. Firstly, in each residual block, we use the commonly adopted cross-attention with noisy latents as the query and audio embeddings $c_{audio}$ as the key and value to compute an attended audio feature. This attended audio feature is then added to the noisy latent features obtained from inter-clip temporal layer, resulting in a new noisy latent features. This provides a preliminary audio condition for the model.
@@
Audio-to-Latents Module. Additionally, as illustrated in Figure 4, we introduce audio-to-latents module to enhance the influence of audio on portrait motion. This module receives various inputs during training, including head movement and expression variances (strongly correlated with portrait motion, details are provided in the Appendix) and audio features (weakly correlated). The module maps these conditions to a shared motion latents space, which replaces the original conditions in subsequent computations. This method allows audio to more accurately influence motion by leveraging the shared motion latents learned from strong conditions, also avoiding focus on irrelevant areas. Specifically, we maintain a set of learnable embeddings. For each input condition, we map it to query features using a fully connected (FC) layer, while the learnable embeddings serve as key and value features for attention computation to obtain a new feature based on the learnable embeddings. This attention computation is conducted on the token dimension. Audio and variance features are transformed into a query feature via different FC layers (token count is 1). For key and value features, the token count corresponds to the number of learnable embedding vectors, which we have set to 128. The FC layers also unify the QKV features to 256 channels for attention computation. The obtained value features, i.e., motion latents, are transformed to 1280 channels via an FC layer and added to the timestep embedding. During training, we sample an input condition for the audio-to-latents module with equal probability from audio embeddings, head absolute movement variance and face expression variance. During testing, we only input audio features to generate motion latents. In fact, expressions and movements can also influence motion generation through motion latents, as demonstrated on the project page.
@@ 3.4 Training Strategies
Conditions Mask and Dropout. In the Loopy framework, various conditions are involved, including the reference image $c_{ref}$, audio features $c_{audio}$, preceding frame motion frames $c_{mf}$, and motion latents representing audio and facial movement conditions.
To better learn the unique information specific to each condition, we use distinct masking strategies for the conditions during the training process. During training, $c_{audio}$ and motion latents are masked to all-zero features with a 10% probability. For $c_{ref}$ and $c_{mf}$, we design specific dropout and mask strategies due to their highly overlapping information. $c_{mf}$ also provides appearance information and is closer to the current clip compared to $c_{ref}$, leading the model to heavily rely on motion frames rather than the reference image.
To address this, $c_{ref}$ has a 15% probability of being dropped, meaning the denoising U-Net will not concatenate features from the reference network during spatial-attention computation. When $c_{ref}$ is dropped, motion frames are also dropped, meaning the denoising U-Net will not concatenate features from the reference network during temporal attention computation. Additionally, motion frames have an independent 40% probability of being masked to all-zero features.
Multistage Training. Following AnimateAnyone (Hu, 2024) and EMO (Tian et al., 2024), we employ a 2-stage training process. In the first stage, the model is trained without temporal layers and the audio condition module. The inputs to the model are the noisy latents of the target single-frame image and reference image latents, focusing the model on an image-level pose variations task. After completing the first stage, we proceed to the second stage, where the model is initialized with the reference network and denoising U-Net from the first stage. We then add the inter/intra- clip temporal modules and the audio condition module for full training to obtain the final model.
@@
where $e_{audio}$ includes all conditions, $e_{ref}$ drops the $c_{audio}$ condition, and $e_{base}$ further drops the $c_{ref}$. The audio ratio is set to 5 and the reference ratio to 3. We use DDIM with 25 denoising steps for inference. Ungenerated parts of motion frames are masked to all-zero features until generated.
@@ 3.5 Experiments
Datasets. For training data, we collected and filtered talking head videos from multiple sources, including face-related datasets, general-purpose datasets, and online platforms, excluding videos with audio that have low lip-sync scores (Chung & Zisserman, 2017), excessive head movement, and rotation. This resulted in 174 hours of training data. The dataset are detailed in the Appendix LABEL:dataset_a. For test sets, we randomly sampled 100 videos from CelebV-HQ (Zhu et al., 2022) (a public high-quality celebrity video dataset with mixed scenes), RAVDESS (Kaggle, ) (a public high-definition indoor talking scene dataset with rich emotions) and HDTF (Zhang et al., 2021). To test the generalization ability of diffusion-based models, we also collected 20 portrait test images, including real people, anime, side face, and humanoid crafts of different materials, along with 20 audio clips, including speeches, singing, rap and emotionally rich speech. We refer to this test set as the openset test set.
@@
Implementation Details. We trained our model using 24 Nvidia A100 GPUs with a batch size of 24, using the AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of 1e-5 to train the model for two stages, each lasting 4 days. The generated video length was set to 12 frames, and the motion frame was set to 124 frames, representing the preceding 124 frames of the current 12-frame video. After temporal segment module, this was abstracted to 20 motion frame latents. During training, the reference image was randomly selected from frames within the video. For the facial motion information required to train audo-to-motion module, we used DWPose (Yang et al., 2023) to detect facial keypoints for the current 12 frames. The variance of the absolute position of the nose tip across these 12 frames was used as the absolute head movement variance. The variance of the displacement of the upper half of the facial keypoints (37 keypoints) relative to the nose tip across these 12 frames was used as the expression variance. The training videos were uniformly processed at 25 FPS and cropped to 512$\times$512 portrait videos.
@@
Metrics and Compared Baselines. We assess image quality using the IQA metric (Wu et al., 2023), video motion stability with VBench’s smooth metric (Huang et al., 2024), and audio-visual synchronization with SyncC and SyncD (Chung & Zisserman, 2017). For the CelebvHQ and RAVDESS test sets, which have corresponding ground truth videos, we also compute FVD (Unterthiner et al., 2019), E-FID (Tian et al., 2024), and FID metrics for comparison. To demonstrate lifelike portrait movement beyond just lip movement, we provide global motion metrics (Glo) and dynamic expression metrics (Exp). DGlo and DExp represent the absolute differences from the ground truth, calculated based on the variance of key points of the nose and upper face, excluding the mouth area. More details are provided in the Appendix. For the openset test set, which lacks ground truth video references, we conducted subjective evaluations. Ten users assessed six dimensions: identity consistency, video synthesis quality, audio-emotion matching, motion diversity, naturalness of motion, and lip-sync accuracy. Participants identified the top-performing method in each dimension. For baseline methods, We compared recent state-of-the-art audio-driven portrait methods based on diffusion models, including Hallo (Xu et al., 2024a), VExpress (Wang et al., 2024) and EchoMimic (Chen et al., 2024), and also included the competitive GAN-based method, SadTalker Zhang et al. (2023b), as a reference. We also included the results of the open-source method Hallo, trained on our dataset according to the official instructions, marked with *.
@@ 3.5.1 Results and Analysis
Performance in Complex Scenarios. CelebV-HQ features videos of celebrities speaking in diverse scenarios, both indoors and outdoors, with various portrait poses. This makes testing on this dataset effectively simulate real-world usage conditions. As shown in Table 1, our method significantly outperforms the other compared methods in most metrics, as evidenced by the comparison videos provided in the supplementary materials. Regarding motion-related metrics, in the dynamic expression metric (Exp), our method closely matches GT, outperforming other methods. For global motion (Glo), our performance is similar to EchoMimic. However, our method distinctly excels in video synthesis quality and lip-sync accuracy. For Hallo trained on our training set (marked with *), the results did not improve, and similar outcomes in subsequent experiments indicate that the design of our method is key to achieving satisfactory results.
@@
SadTalker 2.953 3.843 8.765 171.848 1746.038 36.648 0.554 0.270 0.291 0.368 2.248
Hallo 3.505 4.130 9.079 53.992 742.974 35.961 0.499 0.255 0.301 0.329 2.426
Hallo* 3.467 2.607 10.875 63.142 883.249 48.676 0.489 0.412 0.313 0.420 2.516
VExpress 2.946 3.547 9.415 117.868 1356.510 65.098 0.020 0.166 0.339 0.464 2.414
EchoMimic 3.307 3.136 10.378 54.715 828.966 35.373 2.259 0.640 0.260 0.442 3.018
Loopy 3.780 4.849 8.196 49.153 680.634 33.204 2.233 0.452 0.279 0.309 2.307
@@
SadTalker 3.840 4.304 7.621 22.516 487.924 32.343 0.604 0.120 0.271 0.213 3.270
Hallo 4.393 4.062 8.552 38.471 537.478 19.826 0.194 0.080 0.299 0.243 3.785
Hallo* 4.233 2.878 9.672 47.758 691.103 34.973 0.354 0.129 0.291 0.191 3.621
VExpress 3.690 5.001 7.710 62.388 982.810 26.736 0.007 0.039 0.329 0.283 3.901
EchoMimic 4.504 3.292 9.096 54.115 688.675 21.058 0.641 0.184 0.263 0.182 3.350
Loopy 4.506 4.814 7.798 16.134 394.288 17.017 2.962 0.343 0.260 0.197 3.132
Figure 5: User voting comparisons on the openset test set. The first row presents results for different categories of input images, while the second row shows results for input audio.
Table 3: Comparisons and ablation studies on the openset and HDTF test sets.
Method Openset Test HDTF Test
@@
Loopy (Full) 4.507 6.303 7.749 0.9932 4.017 8.576 6.805 10.443 18.021 1.359
Loopy (HDTF) - - - - 3.894 8.176 7.097 12.742 21.733 1.476
SadTalker 3.749 5.390 9.586 0.9947 3.435 7.320 7.870 24.939 25.353 1.559
EchoMimic 4.447 3.674 10.494 0.9896 3.994 5.546 9.391 18.718 19.015 1.328
Hallo 4.412 5.483 8.798 0.9924 3.922 7.263 7.917 21.717 20.159 1.337
Hallo* 4.413 5.334 8.803 0.9937 3.846 3.579 11.151 29.419 24.689 1.543
VExpress 3.941 4.828 9.572 0.9961 3.482 8.186 7.382 48.038 30.916 1.506
w/o inter-clip temp. 4.335 6.104 8.129 0.9942 3.769 8.355 6.899 11.595 19.277 1.486
w/o TSM 4.386 6.054 8.235 0.9922 3.908 7.920 7.461 11.552 19.023 1.401
w/o A2L 4.428 5.999 8.351 0.9922 3.918 8.127 7.188 10.837 18.153 1.383
1 temp.+ 20 MF 4.072 6.201 8.309 0.9752 3.765 8.494 6.795 11.817 19.819 1.380
$s=1,\ r=2$ 4.461 5.919 8.245 0.9940 3.945 7.848 7.388 14.814 19.665 1.392
$s=2,\ r=2$ 4.453 5.855 8.326 0.9930 3.946 8.172 7.038 13.310 18.690 1.398
$s=3,\ r=2$ 4.443 6.083 8.161 0.9930 3.941 7.985 7.263 11.260 18.655 1.322
$s=4,\ r=1$ 4.424 6.219 8.004 0.9931 3.937 8.419 6.877 10.672 18.210 1.299
mean sample 4.452 5.907 8.199 0.9931 3.865 7.851 7.570 10.506 19.475 1.535
random sample 4.438 6.098 8.144 0.9932 3.865 7.866 7.588 10.824 18.229 1.440
Performance in Emotional Expression. RAVDESS is a high-definition talking scene dataset containing videos with varying emotional intensities. It effectively evaluates the method’s performance in emotional expression. The experimental results are listed in Table 2. As shown by the E-FID (Tian et al., 2024) metric, our method outperforms the compared methods. This is corroborated in the motion dynamics metrics Glo and Exp, where our results are close to the ground truth. Although our lip-sync accuracy is slightly inferior to VExpress, it is important to note that the results generated by VExpress generally lack dynamic motion, as indicated by the Glo, Exp, and E-FID metrics. This static nature can provide an advantage when measured with SyncNet for lip-sync accuracy.
Performance in Openset Scenarios. Compared to objective metrics, subjective evaluations more directly reflect the overall performance of the methods. We compared different input styles (real people, anime, humanoid crafts, and side faces) and various types of audio (speech, singing, rap, and emotional audio) to evaluate the performance of the methods. As shown in Figure 5, Loopy consistently and significantly outperforms the compared methods across these diverse scenarios.
@@ 3.5.2 Ablation Studies
Analysis of Key Components. In this part, we conducted experiments on two datasets: openset and HDTF. The former includes complex and challenging inputs, while the latter features relatively simple scenes. We analyzed the impact of the two key components of Loopy, the inter/intra- clip temporal module and the audio-to-latents module. For the former, we conducted two experiments: (1) Removing the inter-clip temporal layer, similar to previous methods. (2) Removing the temporal segment module and retaining 4 motion frames as in other methods. For the latter, we removed the audio-to-latents module and used only cross-attention for audio feature injection. The results, listed in Table 3, indicate that removing the dual temporal layer design, the temporal segment module and the audio-to-latents module all lead to a decline in overall synthesis quality, with the impact being particularly significant when the dual temporal layer design is removed. The Table 3 also includes experimental results of comparison methods on these two datasets and Loopy’s performance on the HDTF test set when trained without collected data for reference. These results substantiate the efficacy of our proposed modules.
Impact of Long-Term Temporal Dependency. We further investigated the impact of long-term motion dependency on the model and listed the results in Table 3. Initially, we compared the effects of extending the motion frame length to 20 under a single temporal layer setting. While this approach enhanced the model’s dynamics, it significantly degraded overall quality (evidenced by a low smooth value). In contrast, the full model with 20 motion frames ($s$=4, $r$=1) and inter/intra-clip temporal layers improved stability, especially on the openset test set. This indicates that when the motion frame length exceeds the current clip, the dual temporal layer design better utilizes long-term motion information, although improper handling may have negative effects. Regarding long-term motion coverage, we fixed $r$=2 and found that smaller $s$ values performed worse. Combined with other ablation results, this suggests that fewer motion frames contain less motion information, leading the network to extract more appearance information. This validates the effectiveness of long-term motion information in enhancing synthesis results. Additionally, we compared different abstraction strategies for motion frames within the temporal segment module. The default uniform sampling approach proved more effective than average pooling and random sampling, likely because it provides more stable and clearer interval information, aiding the inter-clip temporal layer in learning long-term motion information.